# Unlearning as Ablation: Toward a Falsifiable Benchmark for Generative Scientific Discovery

### v1

@@ Abstract
Bold claims about AI’s role in science—from “AGI will cure all diseases” to promises of radically accelerated discovery—raise a central epistemic question: do large language models (LLMs) truly generate new knowledge, or do they merely remix memorized fragments? We propose unlearning-as-ablation as a falsifiable test of constructive scientific discovery. The method systematically removes a target result and its entire forget-closure (lemmas, paraphrases, and multi-hop entailments) and then evaluates whether the model can re-derive the result from only permitted axioms and tools. Success provides evidence for genuine generative capability; failure exposes current limits. Unlike prevailing motivations for unlearning—privacy, copyright, or safety—our framing repositions it as an epistemic probe for AI-for-Science. We argue that such tests could serve as the next generation of benchmarks, much as ImageNet catalyzed progress in vision: distinguishing models that can merely recall from those that can constructively generate new scientific knowledge. We outline a minimal pilot in mathematics and algorithms, and discuss extensions to physics, chemistry, and biology. Whether models succeed or fail, unlearning-as-ablation provides a principled framework to map the true reach and limits of AI scientific discovery. This is a position paper: we advance a conceptual and methodological argument rather than new empirical results.
@@ 1 Introduction
This framing departs from prevailing motivations for unlearning. Surveys emphasize privacy, copyright, and safety as primary rationales (pmlr-v235-huang24u, ; li2024digitalforgetting, ; liu2025sok, ), with evaluation focused on removal fidelity rather than generative ability. Recent work highlights the difficulty of faithfully removing multi-hop or entangled knowledge (cai2024unraveling, ; li2025faithun, ), while other studies show that forgotten content can often be “relearned” through small finetunes or prompting (zheng2023sam, ). In the safety and compliance setting, these phenomena are treated as risks. In our setting, they define the frontier: as unlearning methods improve in addressing leakage and robustness, the resulting ablations become more faithful, and the corresponding rediscovery benchmarks more stringent. In this way, progress in unlearning directly strengthens our ability to test whether models are capable of constructive scientific generation.
@@ 2.1 Motivations for Unlearning
(1) Privacy and compliance. Regulations such as the General Data Protection Regulation (GDPR) enshrine a “right to be forgotten,” requiring that models support the removal of sensitive or personally identifiable data. Surveys on digital forgetting in LLMs emphasize compliance with privacy law as a central driver of research in this area (li2024digitalforgetting, ; gao2023forgetsurvey, ).
(2) Copyright and intellectual property. LLMs trained on large web scrapes may inadvertently memorize copyrighted text, code, or images. Several works argue that machine unlearning is necessary to respect intellectual property claims and to support takedown requests from rights-holders (liu2025sok, ).
(3) Safety and dual-use knowledge. A third line of work focuses on removing hazardous content: for example, step-by-step instructions for synthesizing explosives or pathogens. Recent benchmarks such as WMDP (zou2023wmdp, ) evaluate whether unlearning can reduce dual-use risks while maintaining general utility. Methods such as CUT (yoon2024cut, ) specifically target hazardous knowledge removal.
@@ 2.2 Methodological Approaches
  • Representation-level interventions: e.g., Amnesic Probing (elazar2021amnesic, ) removes specific linguistic features from hidden states.
@@ 2.3 Evaluation Challenges
  • Entangled knowledge is difficult to erase. Multi-hop unlearning benchmarks show that even if intermediate nodes are removed, models can often reconstruct targets via alternative reasoning chains (cai2024unraveling, ; li2025faithun, ).
  • Suppression vs. removal. SoK papers stress the importance of distinguishing true parameter-level removal from surface-level suppression, where models appear to forget but can be prompted to recall (liu2025sok, ).
  • Relearning and robustness. Empirical work demonstrates that forgotten content can often be “jogged” back into use with minimal finetuning or prompting (zheng2023sam, ).
@@ 3.1 Defining the Forget-Closure
  • Multi-hop reasoning chains where $T$ can be reconstructed indirectly (cai2024unraveling, ; li2025faithun, ).
@@ 3.2 Performing Strong Unlearning
The second step is to apply removal-oriented unlearning across $\mathcal{F}(T)$. Unlike suppression methods that steer generation away from target outputs, removal aims to eliminate relevant information from the parameterization itself. Candidate techniques include gradient-ascent unlearning, targeted fine-tuning, or optimization-based methods evaluated in recent surveys (liu2025sok, ; li2024digitalforgetting, ). To confirm removal, we propose adopting multi-faceted audits:
@@
  • Counterfactual activation probes (inspired by Amnesic Probing) to test whether $T$-related features still reside in hidden states (elazar2021amnesic, ).
  • Robustness tests against “jogging” attacks, where small finetunes or prompting can restore forgotten knowledge (zheng2023sam, ).
@@ 3.4 Why This Matters
This approach connects progress in unlearning directly to progress in measuring scientific discovery. In the safety and compliance literature, challenges such as entanglement, multi-hop reasoning, and relearning are treated as failure modes because they undermine removal fidelity (cai2024unraveling, ; li2025faithun, ; zheng2023sam, ). In our framing, they set the difficulty of the benchmark: the more effectively unlearning methods address these challenges, the more thoroughly the target knowledge is ablated, and the more demanding the rediscovery task becomes. Thus, advances in unlearning translate into sharper tests of whether LLMs truly possess constructive generative capability. Rather than turning flaws into benefits, we highlight that solving these long-standing problems in unlearning is what enables rigorous epistemic evaluation in AI-for-Science.
@@ 4.3 Evaluation Metrics
  • Leakage audits. Performance on paraphrase, multi-hop, and same-answer sets drawn from $\mathcal{F}(T)$, ensuring the model is not recalling forgotten material (cai2024unraveling, ; li2025faithun, ).
  • Utility retention. Accuracy on unrelated benchmarks (e.g., a subset of MMLU) to confirm that unlearning did not degrade general capability (liu2025sok, ).
@@ 5.2 Turning Failure Modes into Probes
Unlearning research has traditionally cast entanglement, multi-hop reasoning, and relearning as obstacles (cai2024unraveling, ; li2025faithun, ; zheng2023sam, ). In our setting, these challenges become useful stress tests. If a model cannot succeed once closure paths are blocked, it indicates that the relevant knowledge was never truly generative. If it can succeed, it demonstrates robustness and constructive capacity. Either way, phenomena previously treated as evaluation headaches become diagnostic instruments for probing the depth of model reasoning.
@@ 5.5 Toward the Next Major Benchmark
A final implication is that unlearning-as-ablation offers a clear path toward the next generation of benchmarks for AI progress. Just as ImageNet catalyzed advances in computer vision by providing a well-defined task on which algorithms could be compared (imagenet2009, ), a benchmark grounded in constructive re-derivation after unlearning could serve as a lodestar for AI-for-Science. Existing evaluations of knowledge regurgitation and short-form reasoning are increasingly saturated—as highlighted by works such as Humanity’s Last Exam (hlx2024, )—suggesting that the next frontier must measure whether models can move beyond retrieval and interpolation to genuine discovery. We believe that such an “unlearning-as-ablation” benchmark could become a distinguishing test of model strength, separating systems that can merely recall from those that can constructively generate new scientific knowledge.

### v2

@@ Abstract
Bold claims about AI’s role in science—from “AGI will cure all diseases” to promises of radically accelerated discovery—raise a central epistemic question: do large language models (LLMs) truly generate new knowledge, or do they merely remix memorized fragments? We propose unlearning-as-ablation as a falsifiable probe of constructive scientific discovery. The idea is to systematically removes a target result together with its forget-closure (supporting lemmas, paraphrases, and multi-hop entailments) and then evaluate whether the model can re-derive the result from only permitted axioms and tools. Success would indicate generative capability beyond recall; failure would expose current limits. Unlike prevailing motivations for unlearning—privacy, copyright, or safety—our framing repositions it as an epistemic probe for AI-for-Science. We outline a minimal pilot in mathematics and algorithms to illustrate feasibility, and sketch how the same approach could later be extended to domains such as physics or chemistry. This is a position paper: our contribution is conceptual and methodological, not empirical. We aim to stimulate discussion on how principled ablation tests could help distinguish models that reconstruct knowledge from those that merely retrieve it, and how such probes might guide the next generation of AI-for-Science benchmarks.
@@ 1 Introduction
This framing departs from prevailing motivations for unlearning. Surveys emphasize privacy, copyright, and safety as primary rationales (Xu et al., 2023, 2024), with evaluation focused on removal fidelity rather than generative ability. Recent work highlights the difficulty of faithfully removing multi-hop or entangled knowledge (Choi et al., 2024; Wang et al., 2025; Shah et al., 2025), while other studies show that forgotten content can often be “relearned” through small finetunes or prompting (Shah et al., 2025). In the safety and compliance setting, these phenomena are treated as risks. In our setting, they define the frontier: as unlearning methods improve in addressing leakage and robustness, the resulting ablations become more faithful, and the corresponding rediscovery benchmarks more stringent. In this way, progress in unlearning directly strengthens our ability to test whether models are capable of constructive scientific generation.
@@ 2.1 Motivations for Unlearning
(1) Privacy and compliance. Regulations such as the General Data Protection Regulation (GDPR) enshrine a “right to be forgotten,” requiring that models support the removal of sensitive or personally identifiable data. Surveys on digital forgetting in LLMs emphasize compliance with privacy law as a central driver of research in this area (Xu et al., 2024).
(2) Copyright and intellectual property. LLMs trained on large web scrapes may inadvertently memorize copyrighted text, code, or images. Several works argue that machine unlearning is necessary to respect intellectual property claims and to support takedown requests from rights-holders (Karamolegkou et al., 2023; Dou et al., 2025; Xu et al., 2024; Ren et al., 2025; Yao et al., 2024).
(3) Safety and dual-use knowledge. A third line of work focuses on removing hazardous content: for example, step-by-step instructions for synthesizing explosives or pathogens. Recent benchmarks such as WMDP (Li et al., 2024) evaluate whether unlearning can reduce dual-use risks while maintaining general utility.
@@ 2.2 Methodological Approaches
  • Representation-level interventions: e.g., Amnesic Probing (Elazar et al., 2021) removes specific linguistic features from hidden states.
@@ 2.3 Evaluation Challenges
  • Entangled knowledge is difficult to erase. Multi-hop unlearning benchmarks show that even if intermediate nodes are removed, models can often reconstruct targets via alternative reasoning chains (Choi et al., 2024; Wang et al., 2025; Shah et al., 2025).
  • Suppression vs. removal. SoK papers stress the importance of distinguishing true parameter-level removal from surface-level suppression, where models appear to forget but can be prompted to recall (Ren et al., 2025).
  • Relearning and robustness. Empirical work demonstrates that forgotten content can often be “jogged” back into use with minimal finetuning or prompting (Lee et al., 2025).
@@ 3.1 Defining the Forget-Closure
  • Multi-hop reasoning chains where $T$ can be reconstructed indirectly (Choi et al., 2024; Wang et al., 2025; Shah et al., 2025).
@@ 3.2 Performing Strong Unlearning
The second step is to apply removal-oriented unlearning across $\mathcal{F}(T)$. Unlike suppression methods that steer generation away from target outputs, removal aims to eliminate relevant information from the parameterization itself. Candidate techniques include gradient-ascent unlearning, targeted fine-tuning, or optimization-based methods evaluated in recent surveys (Ren et al., 2025). To confirm removal, we propose adopting multi-faceted audits:
@@
  • Counterfactual activation probes (inspired by Amnesic Probing) to test whether $T$-related features still reside in hidden states (Elazar et al., 2021).
  • Robustness tests against “jogging” attacks, where small finetunes or prompting can restore forgotten knowledge (Lee et al., 2025).
@@ 3.4 Why This Matters
This approach connects progress in unlearning directly to progress in measuring scientific discovery. In the safety and compliance literature, challenges such as entanglement, multi-hop reasoning, and relearning are treated as failure modes because they undermine removal fidelity (Dou et al., 2025; Choi et al., 2024; Wang et al., 2025; Shah et al., 2025). In our framing, they set the difficulty of the benchmark: the more effectively unlearning methods address these challenges, the more thoroughly the target knowledge is ablated, and the more demanding the rediscovery task becomes. Thus, advances in unlearning translate into sharper tests of whether LLMs truly possess constructive generative capability. Rather than turning flaws into benefits, we highlight that solving these long-standing problems in unlearning is what enables rigorous epistemic evaluation in AI-for-Science.
@@ 4.3 Evaluation Metrics
  • Leakage audits. Performance on paraphrase, multi-hop, and same-answer sets drawn from $\mathcal{F}(T)$, ensuring the model is not recalling forgotten material (Choi et al., 2024; Wang et al., 2025; Shah et al., 2025).
  • Utility retention. Accuracy on unrelated benchmarks (e.g., a subset of MMLU) to confirm that unlearning did not degrade general capability (Ren et al., 2025; Yao et al., 2024).
@@ 5.2 Turning Failure Modes into Probes
Unlearning research has traditionally cast entanglement, multi-hop reasoning, and relearning as obstacles (Choi et al., 2024; Wang et al., 2025; Shah et al., 2025). In our setting, these challenges become useful stress tests. If a model cannot succeed once closure paths are blocked, it indicates that the relevant knowledge was never truly generative. If it can succeed, it demonstrates robustness and constructive capacity. Either way, phenomena previously treated as evaluation headaches become diagnostic instruments for probing the depth of model reasoning.
@@ 5.5 Toward the Next Major Benchmark
A final implication is that unlearning-as-ablation offers a clear path toward the next generation of benchmarks for AI progress. Just as ImageNet catalyzed advances in computer vision by providing a well-defined task on which algorithms could be compared (Deng et al., 2009), a benchmark grounded in constructive re-derivation after unlearning could serve as a lodestar for AI-for-Science. Existing evaluations of knowledge regurgitation and short-form reasoning are increasingly saturated—as highlighted by works such as Humanity’s Last Exam (Phan et al., 2025)—suggesting that the next frontier must measure whether models can move beyond retrieval and interpolation to genuine discovery. We believe that such an “unlearning-as-ablation” benchmark could become a distinguishing test of model strength, separating systems that can merely recall from those that can constructively generate new scientific knowledge.